OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (mathstudent55):

I assume those 3's are exponents, right?

OpenStudy (anonymous):

yes they are

OpenStudy (mathstudent55):

\(x^3 \cdot x^3 \cdot x^3 \cdot x^3 = x^{3 + 3 + 3 + 3} \) The above is correct, but you can certainly simplify it more by adding the exponents on the right side.

OpenStudy (anonymous):

so the final simplified answer would actually be \[x^{12}\] ?

OpenStudy (mathstudent55):

Correct. \(x^3 \cdot x^3 \cdot x^3 \cdot x^3 = x^{3 + 3 + 3 + 3} = x^{12}\) Now let's see the second part of the question. \( x^3 + x^3 + x^3 + x^3 = 4x^3\) In general, \( 4x^3 \ne x^{12}\)

OpenStudy (anonymous):

ohh so the second part of the question would be \[4x^{4}\] right?

OpenStudy (anonymous):

if so thank you very much!

OpenStudy (mathstudent55):

Careful. It's \(4x^3\), not \(4x^4\).

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

my bad, i just typed in the wrong number, but thank you!

OpenStudy (mathstudent55):

wlcm